the density of aluminum is 2.70 g/ml and the density of copper is 8.96 g/ml. what is the mass of copper would occupy the same volume as 5.00g of aluminum?
1
Expert's answer
2019-10-25T05:30:55-0400

Volume of Al:


"V(Al) =\frac{m}{d} = \frac{5}{2.70} = 1.85 ml"

"V(Cu) = V(Al) = 1.85 ml"

"m(Cu) = d\times V = 8.96\times 1.85 = 16.6 g"